# Nearly geodesic surfaces are filling

### v2

@@ 1 Introduction
We summarize Table 1 for convenience.
Properties 22-dim 33-dim
Minimizing measures of a hypersurface minimizes its intersection number with closed geodesics. False in general, only true for geodesic submanifolds (3.2 and 3.22).
Filling geodesic hypersurfaces have some transverse intersections with every geodesic. for hyperbolic $n$-manifolds with $n\geq 3$ (3.18).
Random geodesic hypersurfaces are filling (a conjecture originated from Wu and Xue [WX22]). Open for random Weil-Petersson model (partial results for expander surfaces by [DS23]). in every closed hyperbolic manifold of dim $\geq 3$, for all but finitely many commensurability classes (Theorem 1.1).
There exist infinitely many embedded geodesic submanifolds. No, not even asymptotically geodesic, 1.2.
The support of measure convergence is distinct from that of Hausdorff convergence. for many sequences of geodesics by Thurston ([Thu97, Mar16]). for asymptotically Fuchsian surfaces ([AL25] and Theorem 1.6, [AL25]) unless eventually commensurable.
Table 1: The beautiful analogies and contrasts between the submanifold theory of hyperbolic surfaces and that of hyperbolic $3$-manifolds.
@@ 1.1 The motivations from hyperbolic geometry and random geodesics
The polar coordinate for the $2$-dimensional hyperbolic metric is
(1.3) $g=\,\mathrm{d}r^{2}+\sinh^{2}r\,\mathrm{d}\theta^{2}.$
@@
$\textnormal{Area}(B)<\ell(\partial B).$
Since a genus-$g$ hyperbolic surface $S_{g}$ has an area of $4\pi(g-1)$, its fundamental domain $D\subset\mathbb{H}^{2}$ has a boundary length of at least $4\pi(g-1)$, and thus a filling geodesic must have a length of at least $2\pi(g-1)$. On $S_{g}$, there are only finitely many closed geodesics of length less than $2\pi(g-1)$. Wu and Xue in [WX22, Question on p. 5] ask the following.
@@
As $g\rightarrow\infty$, on a generic closed hyperbolic surface $S$ in the Weil-Petersson model, are most closed geodesics of length significantly greater than $g$ filling?
Dozier and Sapir in [DS23, Section 5.1] construct a sufficiently fine triangulation $\mathcal{D}$ of $S$ and flow boxes centered at the edges of $\mathcal{D}$ so that most sufficiently long geodesics pass through all these boxes and hence are filling. Their ideas apply to hypersurfaces in higher-dimensional manifolds and detect filling. However, for Theorems 1.8 and 1.1, we have adapted an argument different from [DS23] to prove large Fuchsian surfaces are filling by taking advantage of the rigidity of $SL_{2}(\mathbb{R})$-invariant orbits explained below. The crucial difference is that, given any prescribed subset of $\mathcal{G}_{2}M$, all large Fuchsian surfaces must get close and nearly tangent to.
@@ 1.2 The motivations from the rigidity of unipotent flows
There is a huge variety of invariant sets of geodesic flows on the unit tangent bundle $T^{1}S$ of a hyperbolic surface, i.e., dense [KH95, 17.5, Corollary 18.3.5], closed geodesics, geodesic laminations [Mar16, 8.3], and sets with Hausdorff dimension greater than $1$ but less than $3$ [LS18, Theorem 1.3]. The flexibility of such invariant sets is also reflected by the existence of arbitrarily long closed geodesics supported on a proper subsurface of $S$, which are never filling. Although Fuchsian surfaces have principal curvature 0 and geometrically generalize geodesics, it is their connection with dynamic systems that gives rise to the filling topological nature in Theorem 1.8 and Theorem 1.1. Calegari-Marques-Neves [CMN22] initiated the studies of the distribution of minimal surfaces in negatively curved manifolds using homogeneous dynamics and rigidity results.
@@
Let $G$ be a connected Lie group and $U$ a unipotent one-parameter subgroup of $G$. In contrast with geodesic flows, there is a strong rigidity for the orbits of horocycle flow ($U$-orbits): Hedlund’s theorem [Hed36] proves the horocycle flow on a closed hyperbolic surface is minimal, i.e., every horocycle on a closed hyperbolic surface is dense. Shah [Sha91a] classifies the orbit closure of SO(n1,1)SO(n-1,1) for actions on quotients of $SO(n,1)$ for $n\geq 3$; Etienne Ghys pointed out the implications in geometry: the projection of a codimension-one geodesic hyperplane $P\subset\mathbb{H}^{n}$ to a closed hyperbolic $n$-manifold $M$ is either dense or a closed totally geodesic hypersurface. For $n\geq 3$, the stark contrast between SO(n1,1)SO(n-1,1) and geodesic $SO(1,1)$ relies crucially on the behavior of the orbits of nontrivial unipotent $1$-parameter subgroups contained in SO(n1,1)SO(n-1,1), as can be seen in [Sha91a, Section 4-6]. Margulis [Mar89] shows how to effectively study orbit closures in homogeneous spaces of Lie groups by analyzing minimal closed invariant sets of the action of unipotent subgroups. Some other important works include, for example, Dani-Smillie, Dani, Dani-Margulis, which we refer the readers to the references in e.g., [Mar89, Sha91a, Sha91b, Rat91].
Ratner [Rat91] classifies finite ergodic $U$-invariant measures and proves every unipotent subgroup of a connected Lie group $G$ is strictly measure rigid, which establishes Raghunathan’s measure conjecture. Shah proves a theorem [Sha91b, 1.1] which describes geometrically a subset of points whose $U$-orbits are uniformly distributed with respect to a $G$-invariant measure; in particular, combining with the measure classification [Rat91], he in [Sha91b, 1.3] verifies a stronger form of Raghunathan’s orbit closure conjecture for cocompact quotients of $G$. [MS95] implies that a sequence of Fuchsian surfaces $S_{i}$ in distinct commensurability classes becomes equidistributed, i.e., the probability measure induced by the inclusion of $S_{i}$ on the Grassmann bundle $\mathcal{G}_{2}M$ converges to the uniform Liouville measure on $\mathcal{G}_{2}M$, building on [Rat91]. There is a rapid development of effective Ratner’s theorem. Readers are referred to [LM23, LMW22, Yan22] and the references therein.
@@
Lowe [Low21] studies the minimal foliations of Grassmann bundles of negatively curved manifolds. He also proves various quantitative density and measure convergence results for asymptotically Fuchsian surfaces in non-homogeneous metrics, e.g., [Low21, Theorem 1.9]. Jiang [Jia25] generalizes the minimal surface entropy results of [CMN22] to higher-dimensional and noncompact finite-volume hyperbolic manifolds. Thurston in [Thu97, Remark after Corollary 8.8.6] points out that unlike closed hyperbolic $3$-manifolds, a noncompact finite-volume hyperbolic manifold might have infinitely many homotopy classes of $\pi_{1}$-injective surfaces of bounded genus. Jiang [Jia25] proves that the number of homotopy classes of nearly geodesic surfaces of bounded genus remains finite. Lowe [BLM24], following ideas of Labourie [Lab21], Lowe-Neves [LN21], proves that the weak- limit of the probability measures induced by a sequence of asymptotically Fuchsian surfaces is $PSL_{2}(\mathbb{R})$-invariant.
@@
An outline of the paper is the following. In Section 2, we recall some background on quasi-Fuchsian and minimal surfaces, and the metric and measure on Grassmann bundles. In Section 3, we prove Theorems 1.8 and 1.1, 1.2, 1.3, and 1.5. In Section 4, we classify geometric $3$-manifolds containing a filling $\pi_{1}$-injective surface, construct explicit filling surfaces in Euclidean and $\mathbb{H}^{2}\times\mathbb{R}$ geometry, and prove Theorems 1.10 and 1.11.
@@ Acknowledgment
The author would like to thank Wenyuan Yang for suggesting a connection between Theorem 1.8 and the work of Kahn-Marković. He thanks Vladimir Marković for insightful discussions on the asymptotic behaviors of the planes which help formulate Theorem 1.6 and part of Theorem 1.1. He also thanks Nathan M. Dunfield, Mark Hagen, and Zhenghao Rao for many helpful discussions. He thanks his postdoc mentor Yunhui Wu for many helpful discussions in the last few years on hyperbolic geometry and minimal surfaces. Part of the work is completed when the author visits the Institut Fourier. The author would like to thank the institute for its hospitality and thank Andrea Seppi and Zeno Huang for their interest in this work and for helpful discussions.
@@ 2.1 Quasi-Fuchsian surfaces and minimal surfaces
We call $S\subset M$ primitive if $\pi_{1}(S)$ is not a proper subgroup of the fundamental group of another orientable subsurface. Since $M$ is hyperbolic, a Fuchsian surface $S$ of genus $g$ has intrinsic sectional curvature 1-1. The conjugacy classes of surface subgroups $\pi_{1}(S)$ of $\pi_{1}(M)$ are in one-to-one correspondence with the homotopy classes of $\pi_{1}$-injective immersions $f:S\rightarrow M$. We define a local variation of $f(S)$ to be a smooth map
@@
$\frac{d}{dt}\textnormal{Area}(\Psi_{t}(S))|_{t=0}=0$
for every local variation $\Psi$. By [SU82, SY79], a $\pi_{1}$-injective immersion $f:S\rightarrow M$ is homotopic to a smoothly immersed minimal surface. Henceforth, we assume all such immersions $f:S\rightarrow M$ are minimal unless stated otherwise. A $\pi_{1}$-injective surface $S$ is least area if $S$ achieves the smallest area in its homotopy class. For further background on minimal surfaces, see [CM11] and [Jos17, 5.4].
An immersed surface $f:S\hookrightarrow M$ is $(1+\epsilon)$-quasi-Fuchsian if the limit set of a universal cover $\tilde{S}\subset\mathbb{H}^{3}$ is a $(1+\epsilon)$-quasicircle (but not a $(1+\epsilon^{\prime})$-quasicircle for any $\epsilon^{\prime}<\epsilon$). Commensurable surfaces share the same quasi-Fuchsian constants.
@@
If a $(1+\epsilon)$-quasi-Fuchsian minimal surface $S$ satisfies $\lambda_{0}<1$, then
@@
Lemma 2.3 ([Uhl83] and [Sep16]).
Let $M$ be a closed hyperbolic $3$-manifold. There exists an $\epsilon_{2}>0$ (depending only on $M$) such that for $\epsilon<\epsilon_{2}$, every $(1+\epsilon)-quasi-Fuchsian$ surface is homotopic to a unique minimal surface.
Throughout this work, we assume $\epsilon_{0}$ (e.g., in Theorem 1.1) satisfies $\epsilon_{0}\leq\min\{\epsilon_{1},\epsilon_{2}\}$. For recent advances on almost-Fuchsian manifolds and surfaces, see [HLS23] and the references therein.
When analyzing filling properties and geodesic intersections, there is no loss of generality in considering orientable surfaces of minimal genus within each class. On a hyperbolic surface $S$, a primitive closed geodesic $\gamma$ and its cover $\gamma^{n}$ span identical subsets of $S$. Similarly, a least area primitive surface $S$ and other surfaces commensurable to it span identical subsets of $M$, and their preimages in $\mathbb{H}^{3}$ are the same by [FHS83, Theorem 5.3]. A hypersurface of a Riemannian manifold is a closed codimension-$1$ submanifold with the induced Riemannian metric.

### v3

@@ 1 Introduction
We summarize various theorems for hyperbolic $n$-manifolds in Table 1 for convenience.
Properties $n=2$ $n=3$
Minimizing measures of a hypersurface minimizes its intersection number with closed geodesics. False in general, only true for geodesic submanifolds (3.2 and 3.21).
Filling geodesic hypersurfaces have some transverse intersections with every geodesic. for $n\geq 3$ (3.6).
Random geodesic hypersurfaces are filling (Wu and Xue [WX22] asked on a Weil-Petersson random hyperbolic surface of genus $g$, are most closed geodesics of length $\gg g$ filling). Qualitatively, known to experts, e.g. Arana-Herrera [AH21]. Quantitatively, [DS23]: almost every closed geodesic of length $\gg g(\log g)$ is filling. in every closed hyperbolic manifold of dim $\geq 3$, for all but finitely many classes (Theorem 1.6).
There exist infinitely many embedded geodesic hypersurfaces. No, not even asymptotically geodesic, 1.2.
There exists a sequence of hypersurfaces whose measure convergence has distinct support from Hausdorff convergence. for many sequences of geodesics by Thurston ([Thu97, Mar16]). for asymptotically Fuchsian surfaces ([AL25] and Theorem 1.4, [AL25]) unless eventually commensurable.
Table 1: Some analogies and contrasts between the hypersurface theory of hyperbolic surfaces and that of hyperbolic $3$-manifolds.
@@ 1.1 The motivations from hyperbolic geometry and random geodesics
The polar coordinates for the $2$-dimensional hyperbolic metric is
$g=\,\mathrm{d}r^{2}+\sinh^{2}r\,\mathrm{d}\theta^{2}.$ (1.3)
@@
$\textnormal{Area}(B)<\ell(\partial B).$ (1.4)
Since a genus-$g$ hyperbolic surface $S_{g}$ has an area of $4\pi(g-1)$, its fundamental domain $D\subset\mathbb{H}^{2}$ has a boundary length of at least $4\pi(g-1)$. A filling geodesic $\gamma$ cuts the ambient hyperbolic surface into a disjoint union of polygons. Thus by (1.4), a filling geodesic must have a length of at least $2\pi(g-1)$. On $S_{g}$, there are only finitely many closed geodesics of length less than $2\pi(g-1)$. Arana-Herrera [AH21] suggests that all but a quantitatively small number of primitive closed geodesics of a closed, orientable hyperbolic surface are filling. Quantitatively, Wu and Xue [WX22, Question on p. 5] ask the following.
@@
As $g\rightarrow\infty$, on a generic closed hyperbolic surface $S$ in the Weil-Petersson model, are most closed geodesics of length $\gg g$ filling?
Dozier and Sapir in [DS23, Section 5.1] construct a sufficiently fine triangulation $\mathcal{D}$ of $S$ and flow boxes centered at the edges of $\mathcal{D}$ so that most sufficiently long geodesics pass through all these boxes and hence are filling. Their insightful ideas apply to hypersurfaces in higher-dimensional manifolds and detect fillingness. However, for Theorems 1.6 and 1.1, we have adapted an argument different from [DS23] to prove large totally geodesic surfaces are filling by taking advantage of the rigidity of $SL_{2}(\mathbb{R})$-orbits explained below. The crucial difference is that, given any prescribed subset $O$ of $\mathcal{G}_{2}M$, all large totally geodesic surfaces must get close and nearly tangent to $O$.
@@ 1.2 The motivations from the rigidity of unipotent flows
There is a huge variety of invariant sets of geodesic flows on the unit tangent bundle $T^{1}S$ of a hyperbolic surface, i.e., dense [KH95, 17.5, Corollary 18.3.5], closed geodesics, geodesic laminations [Mar16, 8.3], and sets with Hausdorff dimension greater than $1$ but less than $3$ [LS18, Theorem 1.3]. The flexibility of such invariant sets is also reflected by the existence of arbitrarily long closed geodesics supported on a proper subsurface of $S$, which are never filling. Although Fuchsian surfaces have principal curvature 0 and geometrically generalize geodesics, it is their connection with dynamic systems that gives rise to their filling topological nature in Theorem 1.6 and Theorem 1.1.
@@
Let $G$ be a connected Lie group and $U$ a unipotent one-parameter subgroup of $G$. In contrast with geodesic flows, there is a strong rigidity for the orbits of horocycle flow ($U$-orbits): Hedlund’s theorem [Hed36] proves the horocycle flow on a closed hyperbolic surface $S$ is minimal, i.e., every horocycle on $S$ is dense. Let $SO_{0}(n,1)$ denote the connected component of the group of linear transformations on $\mathbb{R}^{n+1}$ preserving the bilinear forms $\langle x,y\rangle=\sum_{i=1}^{n}x_{i}y_{i}-x_{n+1}y_{n+1}.$ Shah [Sha91a] classifies the orbit closure of $SO_{0}(n-1,1)$ for actions on finite-volume quotients of $SO_{0}(n,1)$ (the framed bundle of $\mathbb{H}^{n}$) for $n\geq 3$; Etienne Ghys pointed out the implications in geometry: the projection of a codimension-one geodesic hyperplane $P\subset\mathbb{H}^{n}$ to a closed hyperbolic $n$-manifold $M$ is either dense or a closed totally geodesic hypersurface. For $n\geq 3$, the stark contrast between $SO_{0}(n-1,1)$ and $SO_{0}(1,1)$ (which acts as the geodesic flow) relies crucially on the behavior of the orbits of nontrivial unipotent $1$-parameter subgroups $U$ contained in $SO_{0}(n-1,1)$, as can be seen in [Sha91a, Section 4-6] (the linearization of $U$ is a polynomial, allowing non-divergence estimates which do not hold for geodesic flows). Margulis [Mar89] shows how to effectively study orbit closures in homogeneous spaces of Lie groups by analyzing minimal closed invariant sets of the action of unipotent subgroups. Some other important works include, for example, Dani-Smillie, Dani, Dani-Margulis, which we refer the readers to the references in e.g., [Mar89, Sha91a, Sha91b, Rat91].
Ratner’s celebrated breakthrough [Rat91] classifies finite ergodic $U$-invariant measures and proves every unipotent subgroup $U$ of a connected Lie group $G$ is strictly measure rigid, i.e., for every discrete subgroup $\Gamma\leq G$ and every $U$-invariant Borel probability measure $\mu$ on $\G\Gamma\backslash$ G, the support of $\mu$ is a closed orbit of the subgroup of $G$ preserving $\mu$. Building on [Rat91], [MS95] implies that in a finite-volume hyperbolic $n$-manifold $M$, a sequence of non-commensurable totally geodesic hypersurfaces $S_{i}$ becomes equidistributed, i.e., the probability measure induced by the inclusion of $\mathcal{G}_{n-1}S_{i}$ on the Grassmann bundle $\mathcal{G}_{n-1}M$ converges to the Liouville measure on $\mathcal{G}_{n-1}M$.
@@
Calegari-Marques-Neves [CMN22] initiated the studies of the distribution of minimal surfaces in negatively curved manifolds using homogeneous dynamics and rigidity results. Lowe [Low21] studies the minimal foliations of Grassmann bundles of negatively curved manifolds. He also proves various quantitative density and measure convergence results for asymptotically geodesic surfaces in non-homogeneous metrics, e.g., [Low21, Theorem 1.9]. Jiang [Jia25] generalizes the minimal surface entropy results of [CMN22] to higher-dimensional and noncompact finite-volume hyperbolic manifolds. Thurston in [Thu97, Remark after Corollary 8.8.6] points out that unlike closed hyperbolic $3$-manifolds, a noncompact finite-volume hyperbolic manifold might have infinitely many homotopy classes of $\pi_{1}$-injective surfaces of bounded genus. Jiang [Jia25] proves that the number of homotopy classes of nearly geodesic surfaces of bounded genus remains finite. Lowe [Low21], following ideas of Labourie [Lab21] and Lowe-Neves [LN21], proves that the weak- limit of the probability measures induced by a sequence of asymptotically geodesic surfaces is $PSL_{2}(\mathbb{R})$-invariant. Al Assal [AL25] proves that the space of such limiting measures is exactly the space of $PSL_{2}(\mathbb{R})$-invariant measures.
@@
An outline of the paper is the following. In Section 2, we recall some background on quasi-Fuchsian and minimal surfaces, and the metric and measure on Grassmann bundles. In Section 3, we prove Theorems 1.6 and 1.1, 1.2, 1.3, 1.8, and 1.9. In Section 4, we classify geometric $3$-manifolds containing a filling $\pi_{1}$-injective surface, construct explicit filling surfaces in Euclidean and $\mathbb{H}^{2}\times\mathbb{R}$ geometry, and prove Theorems 1.10 and 1.11.
@@ Acknowledgment
The author would like to thank Wenyuan Yang for suggesting a connection between Theorem 1.6 and the work of Kahn-Marković. He thanks Vladimir Marković for insightful discussions on the asymptotic behaviors of the planes which help formulate Theorem 1.4 and part of Theorem 1.1. He also thanks Nathan M. Dunfield, Mark Hagen, Zhenghao Rao, and Yibo Zhang for many helpful discussions. He thanks his postdoc mentor Yunhui Wu for many helpful discussions in the last few years on hyperbolic geometry and minimal surfaces. The author is truly grateful for anonymous referee’s very long report which significantly improves the paper in many aspects. Part of the work is completed when the author visits the Institut Fourier. The author would like to thank the institute for its hospitality and thank Andrea Seppi and Zeno Huang for their interest in this work and for helpful discussions. The author is partially supported by the start-up grants at Shanghai Institute of Mathematics and Interdisciplinary Sciences and NSFC No. 12501084.
@@ 2.1 Quasi-Fuchsian surfaces and minimal surfaces
A surface $S\hookrightarrow M$ is $\pi_{1}$-injective if the induced homomorphism $\pi_{1}(S)\rightarrow\pi_{1}(M)$ is injective. We call $S\hookrightarrow M$ primitive if $\pi_{1}(S)$ is not a proper subgroup of the fundamental group of another orientable subsurface. Since $M$ is hyperbolic, a Fuchsian surface $S$ of genus $g$ has intrinsic sectional curvature 1-1. The conjugacy classes of surface subgroups $\pi_{1}(S)$ of $\pi_{1}(M)$ are in one-to-one correspondence with the homotopy classes of $\pi_{1}$-injective immersions $f:S\rightarrow M$. We define a local variation of $f(S)$ to be a smooth map
@@
$\frac{\,\mathrm{d}}{\,\mathrm{d}t}\textnormal{Area}(\Psi_{t}(S))|_{t=0}=0$
for every local variation $\Psi$. A $\pi_{1}$-injective surface $S$ is least area if $S$ achieves the smallest area in its homotopy class (which is then necessarily minimal). By [SU82, SY79], a $\pi_{1}$-injective immersion $f:S\rightarrow M$ is homotopic to a smoothly immersed least area surface. Henceforth, we assume all such immersions $f:S\rightarrow M$ are minimal unless stated otherwise. A surface is Fuchsian or totally geodesic if any geodesic on $S$ with respect to its induced Riemannian metric is also a geodesic on $M$. For further background on minimal surfaces, see [CM11] and [Jos17, 5.4].
An immersed surface $f:S\hookrightarrow M$ is $(1+\epsilon)$-quasi-Fuchsian if the limit set of a universal cover $\tilde{S}\subset\mathbb{H}^{3}$ is a $(1+\epsilon)$-quasicircle. Two $\pi_{1}$-injective surfaces $S_{1},S_{2}\subset M$ are commensurable if there exists a $g\in\pi_{1}(M)$ such that $g\pi_{1}(S_{1})g^{-1}\cap\pi_{1}(S_{2})$ is of finite index in both $\pi_{1}(S_{1})$ and $\pi_{1}(S_{2})$. Commensurable surfaces share the same quasi-Fuchsian constants.
@@
If a $(1+\epsilon)$-quasi-Fuchsian minimal surface $S$ satisfies $0<\lambda_{0}<1$, then
@@
Theorem 2.3 ([Uhl83] and [Sep16]).
Let $M$ be a closed hyperbolic $3$-manifold. There exists an $\epsilon_{2}>0$ (depending only on $M$) such that every $(1+\epsilon_{2})-quasi-Fuchsian$ surface is homotopic to a unique minimal surface.
This theorem allows us to speak of the distribution of asymptotically geodesic surfaces. Throughout this work, we assume $\epsilon_{0}$ (e.g., in Theorem 1.1) satisfies $\epsilon_{0}\leq\min\{\epsilon_{1},\epsilon_{2}\}$. For recent advances on almost-Fuchsian manifolds and surfaces, see [HLS23] and the references therein.
When analyzing filling properties and geodesic intersections, there is no loss of generality in considering orientable surfaces of minimal genus within each class modulo homotopy and commensurability. On a hyperbolic surface $S$, a primitive closed geodesic $\gamma$ and its cover $\gamma^{n}$ span identical subsets of $S$. Similarly, it follows from the uniqueness of almost Fuchsian minimal surface in its homotopy class that a least area primitive almost Fuchsian surface $S$ and commensurable least area surfaces span identical subsets of $M$, and their preimages in $\mathbb{H}^{3}$ are the same by [FHS83, Theorem 5.3]. A hypersurface of a Riemannian manifold is a closed codimension-$1$ submanifold with the induced Riemannian metric.